# An Optimism-based Approach to Online Evaluation of Generative Models

### v1

@@ 7.1 Empirical Results of Online FID-based Evaluation
Performance on pretrained generators under different embeddings. The results are summarized in Figure 1, where each row and each column correspond to one dataset and one performance metric, respectively. The results show that FID-UCB (blue) outperforms naive-UCB (orange) and the greedy algorithm (green). We observe that naive-UCB fails to distinguish between different generators and makes uniform selections on the CIFAR10 and the ImageNet datasets, while FID-UCB recognizes the optimal one. This shows that the generator-dependent optimistic FID (6) can better exploit the properties of the generator, which is key to the sample-efficient online evaluation. On the FFHQ dataset, all the generators attain very similar FID scores (the largest gap is less than 0.7), which may explain why FID-UCB has a similar performance to naive-UCB. We reran these experiments using DINOv2-ViT-L/14-based [33] and CLIP-based [7] embeddings. (The results are summarized in Figures 5 and 6 in Appendix C, respective.) The numerical results show that FID-UCB can maintain satisfactory performance over different embeddings.
CIFAR10 (G=5)
ImageNet (G=5)
[…]
Figure 1: Performance of FID-UCB, naive-UCB, and the greedy algorithm on various datasets. The embeddings of the images are extracted by the pretrained InceptionNet.V3.
Performance on variance-controlled generators. We summarize the results in Figure 2. FID-UCB (blue) significantly outperforms naive-UCB and the greedy algorithm. The results for DINOv2-based and CLIP-based FID are provided in Figures 5 and 6 in Appendix C. We ran the evaluation on AFHQ Dog dataset, and the results are presented in Figure 7 in the Appendix. Furthermore, we display the generated images to visualize the quality and diversity of the samples queried by different algorithms. The results suggest that FID-UCB consistently queries images that are diverse and of high quality. In contrast, naive-UCB kept querying a proportion of images from the collapsed generator (the one with the smallest truncation parameter), and the greedy algorithm even stuck to it.
FFHQ (G=20)
Figure 2: Performance of FID-UCB, naive-UCB, and the greedy algorithm for evaluating the variance-controlled generators on the FFHQ dataset.
Step 500
[…]
7.2 Results of Online IS-Based Evaluation
@@
CIFAR10 (G=5)
Figure 3: Performance of IS-UCB, naive-UCB, and the greedy algorithm on various datasets.
ImageNet (G=5) [Uncaptioned image]
[…]
Performance on variance-controlled generators. The results are summarized in Figure 4. Generated samples are plotted in Figure 8 in the Appendix. The results on the AFHQ Dog dataset are presented in Figure 9 in the Appendix.
FFHQ (G=20)
Figure 4: Performance of IS-UCB, naive-UCB, and the greedy algorithm for evaluating the variance-controlled generators on the FFHQ dataset.
@@ 8 Conclusion
In this work, we studied an online learning problem aiming to identify the generative model with the best evaluation score among a set of models. We proposed the optimism-based FID-UCB and IS-UCB algorithms to address the online learning task and showed satisfactory numerical results for their application to image-based generative models. An interesting future direction to our work is to explore the application of other online learning frameworks, such as Thompson sampling, to address the online evaluation problem. Also, while we mostly focused on standard image datasets for testing the proposed algorithms, the application of the proposed FID-UCB and IS-UCB to text-based and video-based generative models will be interesting for future studies. Such an application only requires applying text and video-based embeddings, as in Video IS [35], FVD [42], and FTD [19].
Appendix A Proofs in Section 5: FID Evaluation
A.1 Proof of Theorem 5.1: Optimistic FID Score
@@ Appendix A Proofs in Section 5: FID Evaluation
The proof of Theorem 5.1 is based on Theorem A.1 and Lemma A.2 in the Appendix. Specifically, Inequality (16) in Theorem A.1 in Appendix A.3 decomposes the estimation error of the empirical FID score (4) into the errors in estimating the mean (i.e., $\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}$) and the covariance matrix (i.e., $\|\widehat{\Sigma}_{g}^{n}-\Sigma_{g}\|_{2}$). Next, Lemma A.2 in Appendix A.4 derives generator-dependent concentration errors for the mean and the covariance matrix. Combining ∎
A.2 Proof of Theorem 5.2: Regret of FID-UCB
@@
Recall that $g_{t}$ denotes the generator picked at the $t$-th step. For convenience, we denote by $\widehat{\textup{FID}}_{g_{t}}$ the optimistic FID of generator $g_{t}$ computed at step $t$. We denote by $n_{t}$ the number of images generated by model $g_{t}$ after $t$ steps. First, by Theorem A.2 and a union bound over $T$ steps, with probability at least $1-\delta$, we have that $\textup{FID}_{g}\geq\widehat{\textup{FID}}_{g}$ for any step $t\in[T]$ and $g\in[G]$. Hence, we have that
$\textup{Regret}(T)=\sum_{t=1}^{T}\left(\textup{FID}_{g_{t}}-\textup{FID}^{*}\right)\leq\sum_{t=1}^{T}\left(\textup{FID}_{g_{t}}-\widehat{\textup{FID}}_{g_{t}}\right).$
@@
By the definition of $\widehat{\textup{FID}}_{g_{t}}$, we further derive that
$\textup{Regret}(T)\leq$ $\sum_{t=1}^{T}\left(\textup{FID}_{g_{t}}-\widetilde{\textup{FID}}_{g_{t}}^{n_{t}}+{\mathcal{B}}^{n_{t}}_{g_{t}}\right)\leq 2\sum_{t=1}^{T}{\mathcal{B}}^{n_{t}}_{g_{t}},$
where ${\mathcal{B}}^{n}_{g}$ is the bonus function given by Equation (7). Therefore, the regret is further bounded by
@@
$\widetilde{O}\left(\max_{g}\left\{G\cdot{\mathcal{I}}_{g}\cdot\log(T),\sqrt{G({\mathcal{I}}_{g}+\textup{Tr}[\Sigma_{g}])\cdot T},(\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\cdot\sigma_{g}\|\Sigma_{g}\|_{2}^{\frac{1}{2}}\bm{r}^{\frac{1}{4}}(\Sigma_{g}))\cdot G^{\frac{1}{4}}T^{\frac{3}{4}}\right\}\right),$
where logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$. Here, we use the fact that
$\sum_{t=1}^{T}\frac{1}{n_{t}^{\alpha}}\leq\frac{(2b)^{\alpha}}{1-\alpha}\sum_{g=1}^{G}(n_{g}(T))^{1-\alpha}\leq\frac{(2b)^{\alpha}}{1-\alpha}G^{\alpha}T^{1-\alpha}\text{, where }0\leq\alpha<1,$
$\sum_{t=1}^{T}\frac{1}{n_{t}}\leq 2b\sum_{g=1}^{G}\log(n_{g}(T))\leq 2bG\log(T),$
where $b$ is the batch size, and $n_{g}(T)$ is the number of images generated by model $g$ at the last step $T$. ∎
A.3 Concentration of Empirical FID (4)
Theorem A.1 (Concentration of empirical FID (4)).
Assume the covariance matrix $\Sigma_{\textup{r}}\succ\bm{0}$ is positive strictly definite, and we denote by $\Sigma_{\textup{r}}=Q_{\textup{r}}^{\top}\Lambda_{\textup{r}}Q_{\textup{r}}$ its eigendecomposition. Then, with probability at least $1-\frac{\delta}{3}$, we have that
@@ A.3 Concentration of Empirical FID (4)
$\left|\widetilde{\textup{FID}}_{g}^{n}-\textup{FID}_{g}\right|\leq$ $\underbrace{(\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}+2\|\mu_{g}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}}_{\text{(1) error of $\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$}}+\underbrace{\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}}_{\text{(2) error of $2\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$}}$ (16)
@@
$+\underbrace{\textup{Tr}(\Sigma_{g})\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{2}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}\right)}_{\text{(3) error of $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$}}.$

### v3

@@ 7.1 Results of Online FD-based Evaluation and Selection
Performance on different embeddings. The results for CLIP embeddings are summarized in Figure 2, where each column and each row correspond to one dataset and one performance metric, respectively. The results show that FD-UCB outperforms Naive-UCB and the Greedy algorithm. We observe that naive-UCB converges to the best model much slower than FD-UCB, which suggests that the generator-dependent and data-driven bonus function can better exploit the properties of the generator, which is key to the sample-efficient online evaluation. We also test the InceptionNet.V3 and DINOv2 embeddings, and the results are summarized in Figures 7 and 8 in the Appendix. FD-UCB can maintain satisfactory performance over different embeddings.
Performance on variance-controlled generators. We summarize the results on the FFHQ and AFHQ-Dog datasets in Figures 910, and 11 in the Appendix. FD-UCB consistently outperformed the baselines.
Performance on video and audio data. We consider the Fréchet Video Distance (Unterthiner et al.,, 2019, FVD) and Fréchet Audio Distance (Kilgour et al.,, 2019, FAD) metrics for video and audio generation, respectively. We synthesize four arms (generators) utilizing the MSR-VTT (Xu et al.,, 2016) and Magnatagatune (Law et al.,, 2009) datasets. When an arm is selected, a video/audio clip is sampled from the dataset and perturbed by Gaussian noises with an arm-specific probability. Results on the audio data are summarized in Figure 6. Additional results can be found in Figure 12 in the Appendix.
Figure 6: Online FAD-based evaluation and selection among four synthetic audio data. The embeddings are extracted by VGGish (Hershey et al.,, 2017). Results for the OPR metric are reported and averaged over 20 trials.
7.2 Results of Online IS-Based Evaluation and Selection
@@
Performance on variance-controlled generators. The results are summarized in Figure 4. The results suggest that FD-UCB consistently queries images that are diverse and of high quality. In contrast, Naive-UCB and the Greedy algorithm kept querying a proportion of images from the collapsed generators with smaller truncation parameters. The AFHQ-Dog dataset results are in the Appendix, Figure 13.
@@ 8 CONCLUSION
In this work, we studied an online learning problem aiming to identify the generative model with the best evaluation score among a set of models. We proposed the multi-armed bandit-based FD-UCB and IS-UCB algorithms to address the online learning task and showed satisfactory numerical results for their application to image-based generative models. An interesting future direction to our work is to explore the application of other online learning frameworks, such as Thompson sampling, to address the online evaluation problem. Also, while we mostly focused on standard image datasets for testing the proposed algorithms, the application of the proposed FD-UCB and IS-UCB to real text-based and video-based generative models will be interesting for future studies. Such an application only requires applying text and video-based embeddings, as in Video IS (Saito et al.,, 2020), FVD (Unterthiner et al.,, 2019), and FBD (Xiang et al.,, 2021). Also, the extension of our MAB framework to a contextual bandit algorithm for prompt-guided generative models, as studied by Hu et al., (2024), will remain for future studies.
Acknowledgments
The work of Farzan Farnia is partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China, Project 14209920, and is partially supported by CUHK Direct Research Grants with CUHK Project No. 4055164 and 4937054. The authors would also like to thank the anonymous reviewers for their constructive feedback and suggestions.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Yes]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes]
    4. (d)
      Information about consent from data providers/curators. [Yes]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Appendix A PROOFS IN SECTION LABEL:SEC:5: FD-BASED EVALUATION
A.1 Proof of Theorem 1: Optimistic FD Score
@@
The proof of Theorem 1 is based on Theorem 4 and Lemma 1 in the Appendix. Specifically, Inequality (18) in Theorem 4 in Appendix A.3 decomposes the estimation error of the empirical FD score (4) into the errors in estimating the mean (i.e., $\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}$) and the covariance matrix (i.e., $\|\widehat{\Sigma}_{g}^{n}-\Sigma_{g}\|_{2}$): With probability at least $1-\frac{\delta}{3}$, it holds that
$\left|\widetilde{\textup{FD}}_{g}^{n}-\textup{FD}_{g}\right|\leq$ $2\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right)+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}$
$+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right).$
Next, Lemma 1 in Appendix A.4 derives generator-dependent concentration errors for the mean and the covariance matrix: With probability at least $1-\frac{2\delta}{3}$, it holds that
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}\leq\sqrt{\frac{1}{n}\left(\sqrt{8\cdot\textup{Tr}[\Sigma_{g}^{2}]\log\left(\frac{6}{\delta}\right)}+8\|\Sigma_{g}\|_{2}\log\left(\frac{6}{\delta}\right)\right)}=:\Delta_{\mu_{g}}^{n}$
$\left\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\right\|_{2}\leq 20\kappa^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+(\Delta_{\mu_{g}}^{n})^{2}=:\Delta_{\Sigma_{g}}^{n}$
Therefore, let
${\mathcal{B}}_{g}^{n}=2\Delta_{\mu_{g}}^{n}\cdot\left(\Delta_{\mu_{g}}^{n}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right)+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\Delta_{\Sigma_{g}}^{n}}+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right),$
and we have that $\widetilde{\textup{FD}}_{g}^{n}-{\mathcal{B}}_{g}^{n}\leq\textup{FD}_{g}$ holds with probability at least $1-\delta$, which concludes the proof. ∎
A.2 Regret of FD-UCB
Theorem 3 (Regret of FD-UCB).
Under the same conditions in Theorem 1, with probability at least $1-\delta$, the regret of the FD-UCB algorithm with a batch size $b\in{\mathbb{N}}_{+}$ is bounded by
$\widetilde{O}\left(\frac{D_{1}^{\mathcal{G}}\cdot G^{1/4}}{b^{1/4}}T^{3/4}+\frac{D_{2}^{\mathcal{G}}\cdot G^{1/2}}{b^{1/2}}T^{1/2}+\frac{D_{3}^{\mathcal{G}}\cdot G}{b}\log T+N\cdot G\right),$ (14)
after running on $G:=|{\mathcal{G}}|$ models for $T\in{\mathbb{N}}_{+}$ steps, where $D_{1}^{\mathcal{G}}$ and $D_{2}^{\mathcal{G}}$ are model-dependent parameters defined in Equations (15)-(17), and the logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$.
@@ A.2 Regret of FD-UCB
Recall that $g_{t}$ denotes the generator picked at the $t$-th step. For convenience, we denote by $\widehat{\textup{FD}}_{g_{t}}$ the optimistic FD of generator $g_{t}$ computed at step $t$. We denote by $n_{t}$ the number of images generated by model $g_{t}$ at the beginning of the $t$-th step. First, by Theorem 1 and a union bound over $T$ steps, with probability at least $1-\delta$, we have that $\textup{FD}_{g_{t}}\geq\widehat{\textup{FD}}_{g_{t}}$ for any step $t\in[T]$ and $g\in[G]$. Hence, we have that
$\textup{Regret}(T)=O(N\cdot G)+\sum_{t=1}^{T}\left(\textup{FD}_{g_{t}}-\textup{FD}^{*}\right)\leq O(N\cdot G)+\sum_{t=1}^{T}\left(\textup{FD}_{g_{t}}-\widehat{\textup{FD}}_{g_{t}}\right),$
@@
where $\textup{FD}^{*}=\min_{g}\textup{FD}_{g}$, and the term $O(N\cdot G)$ corresponds to the regret incurred in the burn-in sampling phase. Further, by the definition of $\widehat{\textup{FD}}_{g_{t}}$, we further derive that
$\textup{Regret}(T)\leq$ $O(N\cdot G)+\sum_{t=1}^{T}\left(\textup{FD}_{g_{t}}-\widetilde{\textup{FD}}_{g_{t}}^{n_{t}}+{\mathcal{B}}^{n_{t}}_{g_{t}}\right)$
$\leq$ $O(N\cdot G)+2\sum_{t=1}^{T}{\mathcal{B}}^{n_{t}}_{g_{t}}$
@@
$\leq$ $\widetilde{O}\left(N\cdot G+\sum_{t=1}^{T}\left(\Delta_{\mu_{g_{t}}}^{n_{t}}\cdot\left(\Delta_{\mu_{g_{t}}}^{n_{t}}+\|\mu_{g_{t}}-\mu_{\textup{r}}\|_{2}\right)+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{\Delta_{\Sigma_{g_{t}}}^{n_{t}}}+\textup{Tr}[\Sigma_{g_{t}}]\sqrt{\frac{1}{n_{t}}}+\frac{\|\Sigma_{g_{t}}\|_{2}}{n_{t}}\right)\right),$
where we utilize $\|\widehat{\mu}_{g_{t}}^{n_{t}}-\mu_{\textup{r}}\|_{2}\leq\Delta_{\mu_{g_{t}}}^{n_{t}}+\|\mu_{g_{t}}-\mu_{\textup{r}}\|_{2}$ in the last inequality. Note that
@@
$\Delta_{\mu_{g_{t}}}^{n_{t}}=\sqrt{\frac{1}{n_{t}}\left(\sqrt{8\cdot\textup{Tr}[\Sigma_{g_{t}}^{2}]\log\left(\frac{6T}{\delta}\right)}+8\|\Sigma_{g_{t}}\|_{2}\log\left(\frac{6T}{\delta}\right)\right)}=\widetilde{O}\left(n_{t}^{-1/2}\sqrt{\sqrt{\textup{Tr}[\Sigma_{g_{t}}^{2}]}+\|\Sigma_{g_{t}}\|_{2}}\right)$
and
$\Delta_{\Sigma_{g_{t}}}^{n_{t}}=$ $20\kappa^{2}\|\Sigma_{g_{t}}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g_{t}})+\log(3T/\delta)}{n_{t}}}+(\Delta_{\mu_{g_{t}}}^{n_{t}})^{2}$
$=$ $\widetilde{O}\left(n_{t}^{-1/2}\sqrt{\|\Sigma_{g_{t}}\|_{2}^{2}(\bm{r}(\Sigma_{g_{t}})+1)}+n_{t}^{-1}\left(\sqrt{\textup{Tr}[\Sigma_{g_{t}}^{2}]}+\|\Sigma_{g_{t}}\|_{2}\right)\right).$
Hence, the regret is further bounded by
@@
$\widetilde{O}\Bigg{(}N\cdot G+$ $\sum_{t=1}^{T}n_{t}^{-1/4}\cdot\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\left(\|\Sigma_{g_{t}}\|_{2}^{2}(\bm{r}(\Sigma_{g_{t}})+1)\right)^{1/4}$
$+n_{t}^{-1/2}\cdot\left((\|\mu_{g_{t}}-\mu_{\textup{r}}\|_{2}+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}])\left(\sqrt{\textup{Tr}[\Sigma_{g_{t}}^{2}]}+\|\Sigma_{g_{t}}\|_{2}\right)^{1/2}+\textup{Tr}[\Sigma_{g_{t}}]\right)$
$+n_{t}^{-1}\cdot\left(\sqrt{\textup{Tr}[\Sigma_{g_{t}}^{2}]}+\|\Sigma_{g_{t}}\|_{2}\right)\Bigg{)}.$
Let
$D_{1}^{\mathcal{G}}:=\max_{g}\left\{\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\left(\|\Sigma_{g}\|_{2}^{2}(\bm{r}(\Sigma_{g})+1)\right)^{1/4}\right\},$ (15)
$D_{2}^{\mathcal{G}}:=\max_{g}\left\{(\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}])\left(\sqrt{\textup{Tr}[\Sigma_{g}^{2}]}+\|\Sigma_{g}\|_{2}\right)^{1/2}+\textup{Tr}[\Sigma_{g}]\right\},$ (16)
$D_{3}^{\mathcal{G}}:=\max_{g}\left\{\sqrt{\textup{Tr}[\Sigma_{g}^{2}]}+\|\Sigma_{g}\|_{2}\right\}.$ (17)
Then, we have that
$\textup{Regret}(T)\leq\widetilde{O}\left(\frac{D_{1}^{\mathcal{G}}\cdot G^{1/4}}{b^{1/4}}T^{3/4}+\frac{D_{2}^{\mathcal{G}}\cdot G^{1/2}}{b^{1/2}}T^{1/2}+\frac{D_{3}^{\mathcal{G}}\cdot G}{b}\log T+N\cdot G\right)$
where we use the fact that
$\sum_{t=1}^{T}\frac{1}{n_{t}^{\alpha}}\leq\frac{2^{\alpha}}{b^{\alpha}(1-\alpha)}\sum_{g=1}^{G}(N_{g}(T))^{1-\alpha}\leq\frac{2^{\alpha}}{b^{\alpha}(1-\alpha)}G^{\alpha}T^{1-\alpha}\text{, where }0\leq\alpha<1,$
$\sum_{t=1}^{T}\frac{1}{n_{t}}\leq\frac{2}{b}\sum_{g=1}^{G}\log(N_{g}(T))\leq\frac{2G}{b}\log T,$
where $b$ is the batch size, and $N_{g}(T)$ is the number of picks of model $g$ at the last step $T$. Therefore, we conclude the proof. ∎
A.3 Concentration of Empirical FD (4)
The following theorem expresses the estimation error of the empirical FD by the concentration of the sample mean and covariance matrix, which facilitates the derivation of the bonus function.
Theorem 4 (Concentration of empirical FD (4)).
Assume the covariance matrix $\Sigma_{\textup{r}}\succ\bm{0}$ is positive strictly definite. Then, with probability at least $1-\frac{\delta}{3}$, we have that
$\left|\widetilde{\textup{FD}}_{g}^{n}-\textup{FD}_{g}\right|\leq$ $2\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right)+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}$ (18)
$+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6}{\delta}\right)}+\frac{8\|\Sigma_{g}\|_{2}}{n}\log\left(\frac{6}{\delta}\right).$